OpenStudy (anonymous):

\[f:\mathbb{R} \rightarrow \mathbb{R} \] is continious with \[Im(f) \subset \mathbb{Z}\] then \[f\] is constant. Give a short reason why or give an example why its not.

hartnn (hartnn):

@jim_thompson5910 @zepdrix any ideas ?

OpenStudy (anonymous):

if the language is not clear enough tell me, so i can try to make it more preciser, my english is not very well in some cases its possible i use wrong words, for example i am suspecious about constant in german its konstant hope has same meaning in english :)

hartnn (hartnn):

@Directrix @cwrw238

hartnn (hartnn):

@oldrin_batuku @satellite73 @amistre64

hartnn (hartnn):

@SithsAndGiggles , any ideas ?

hartnn (hartnn):

@mukushla !!

OpenStudy (anonymous):

Hello what is \(Im(f)\) ?

OpenStudy (anonymous):

Im is Image

OpenStudy (anonymous):

never heard of it... what math u are in?

OpenStudy (anonymous):

calculus

OpenStudy (anonymous):

sorry not image, Im = imaginary

OpenStudy (anonymous):

it must be image, same as range...it cant be imaginary because domain and range of \(f\) are made of real numbers according to problem statement

OpenStudy (anonymous):

ok..

OpenStudy (nincompoop):

R -> R real numbers

OpenStudy (anonymous):

i found the definition For z = x + iy means x Real part, y is the imaginary part of z. Notation: Re(z) = Im(z)

OpenStudy (bahrom7893):

Re(z) = x, Im(z) = y. Let me see if I can dig up my old notes.

OpenStudy (bahrom7893):

http://www.mediafire.com/download/ixj572fp5leh5lm/Complex_Variables_Notes_-_1.PDF If anyone wants them for future reference, enjoy :)

OpenStudy (anonymous):

considering \(Im\) as image of \(f\) for proving use the fact that \(\mathbb{Z}\) is a Discontinuous subset of \(\mathbb{R}\)

OpenStudy (cruffo):

Can we just use the definition of continuity to show that it's can't jump from one integer value to the next, basically just what @mukushla said

OpenStudy (cruffo):

What would the limit at infinity be?